find the equation of cone whose vertex is(5, 4,3) and base curve 3x2+2y2=6, y+z=0.
1
Expert's answer
2021-07-15T04:29:02-0400

The vertex "V(5,4,3)". Let "a,b,c" be the direction ratios of generator of the cone.

Then the equations of the generator are


"\frac{x-5}{a}=\frac{y-4}{b}=\frac{z-3}{c}=t"

The coordinates of any points on the generator are "(5+at,4+bt,3+ct)".

For some "t\in \R, (5+at,4+bt,3+ct)" lies on the guiding curve.

Therefore "3(5+at)^2+2(4+bt)^2=6" and "4+bt+3+ct=0" .

Thus, "t=-\dfrac{7}{b+c}".

From this we get


"3(5-\dfrac{7a}{b+c})^2+2(4-\dfrac{7b}{b+c})^2=6"

"3(5b+5c-7a)^2+2(4b+4c-7b)^2=6(b+c)^2"

"75b^2+75c^2+147a^2+150bc-210ab-210ac"

"+18b^2-48bc+32c^2-6b^2-12bc-6c^2=0"

"147a^2+87b^2+101c^2-210ab-210ac+90bc=0"

"147(x-5)^2+87(y-4)^2+101(z-3)^2"

"-210(x-5)(y-4)-210(x-5)(z-3)"

"+90(y-4)(z-3)=0"

"147x^2-1470x+3675+87y^2-696y+1392"

"+101z^2-606z+909-210xy+840x+1050y"

"-4200-210xz+630x+1050z-3150+90yz"

"-270y-360z+1080=0"

The equation of cone is


"147x^2+87y^2+101z^2-210xy-210xz+90yz"

"+84y+84z-294=0"